OpenStudy (anonymous):

According to the synthetic division below, which of the following statements are true? Check all that apply. 2(3 -11 10) 6 -10 ____________ 3 -5 0 A. (3x2 - 11x + 10) (x - 2) = (3x - 5) B. (3x2 - 11x + 10) (x + 2) = (3x - 5) C. (x - 2) is a factor of 3x2 - 11x + 10. D. (x + 2) is a factor of 3x2 - 11x + 10. E. The number -2 is a root of F(x) = 3x2 - 11x + 10. F. The number 2 is a root of F(x) = 3x2 - 11x + 10.

OpenStudy (freckles):

\[\frac{P(x)}{x-c}=Q(x)+\frac{R}{x-c} \\ \\ \text{ where } Q(x) \text{ is quotient } \\ \text{ and } R \text{ is remainder } \\ \text{ if } R=0 \text{ then } (x-c) \text{ divides } P(x) \\ \text{ and both } (x-c) \text{ and } Q(x) \text{ are factors of } P(x) \\ \text{ and so } P(x)=Q(x)(x-c) \\ \text{ also } x=c \text{ is a root since when } x=c \text{ we have } P(c)=Q(c)(c-c)=0 \\ \]

OpenStudy (marcelie):

would that be a?

OpenStudy (anonymous):

your goona have to dumb that down for me

OpenStudy (freckles):

the remainder is the number left over in the division

OpenStudy (freckles):

what number is leftover in your division above

OpenStudy (anonymous):

would'nt it be 3, -5, 0

OpenStudy (freckles):

no

OpenStudy (freckles):

the remainder is not 3 different numbers

OpenStudy (freckles):

the number left over is the last number in the last column

OpenStudy (anonymous):

so 0

OpenStudy (freckles):

yep so we do have R=0 that is the remainder is 0 just apply the other things I mentioned above

OpenStudy (freckles):

can you identify what P(x) is and what c is

OpenStudy (freckles):

and what Q(x) is

OpenStudy (anonymous):

so the answers for the questions would be A, C, and F

OpenStudy (freckles):

A is not correct unless you made a type-o in A

OpenStudy (anonymous):

ok not A what about the rest

OpenStudy (freckles):

so there isn't a type-o in A? you wrote A exactly as it was written in your worksheet or whatever you are getting these questions from?

OpenStudy (freckles):

(3x^2-11x+10)/(x-2)=3x-5 would have been right but (3x^2-11x+10)(x-2)=3x-5 is totally not true

OpenStudy (freckles):

\[\frac{P(x)}{x-c}=Q(x)+\frac{R}{x-c}\]  so you are right C and F are definitely true but I'm still concerned you have mistyped A

OpenStudy (anonymous):

so would b be right

OpenStudy (freckles):

no a 2nd deg polynomial times a 1st degree polynomial is a 3rd deg polynomial not a first degree polynomial did you read what I said? I think there is a type-0 in A I showed you by changing the operation from times to division would make A correct

OpenStudy (freckles):

for example x^2*x is x^3 this would not magically end up being x

OpenStudy (freckles):

unless you had x^2/x which is x

OpenStudy (freckles):

did you read what I typed earlier which was this: "(3x^2-11x+10)/(x-2)=3x-5 would have been right but (3x^2-11x+10)(x-2)=3x-5 is totally not true "

OpenStudy (anonymous):

i checked it there was a typo sorry A. (3x2 - 11x + 10) / (x - 2) = (3x - 5)

OpenStudy (freckles):

ok

OpenStudy (freckles):

so you do have 3 answers then